OpenStudy (anonymous):

a/6 - 10 = a/3 An equivalent form of this equation is a - 10 = 2a a - 60 = 6a a - 60 = 2a

OpenStudy (anonymous):

The last one C Explanation: A/6 - 10 = A/3 Your initial reaction should be to eliminate the fractions. You can accomplish this by multiplying the equation by the Lowest Common Denominator (or LCD). So, ask yourself what is the lowest common multiple of 3 and 6 (because 3 and 6 are your denominators)? You should come up with 6. So, multiply the equation by 6. 6(A/6 - 10 = A/3) Now you have A - 60 = 2A because you have to multiply the 6 to every term. 6 * a/6 cancels out the 6 in the denominator leaving you with just A, -10 * 6 equals -60, and 6 * A/3 equals 2A because 6/'3=2. A - 60 = 2A So, the answer is C.

OpenStudy (anonymous):

Thanks so much!:)

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

http://www.cymath.com/answer.php?q=a%2F6-10%3Da%2F3

OpenStudy (anonymous):

@Smileyside Mathway.com is an amazing website for math help! You should check it out.

OpenStudy (anonymous):

i have that site already in my favorites

OpenStudy (anonymous):

It's pretty awesome, isn't it? I absolutely love it!

OpenStudy (anonymous):

yea...

OpenStudy (anonymous):

i thought i blocked you, it says i blocked you

OpenStudy (anonymous):

I blocked you yesterday. But that's in the past and I am sorry about that! Can we be friends instead?

OpenStudy (anonymous):

sure... but it is just a site so no fighting or i will get a mod

OpenStudy (anonymous):

...

OpenStudy (anonymous):

That's why I just apologized and asked if we could be friends..lol

OpenStudy (anonymous):

ok lol.....